Coefficient of transmission and coefficient of reflection for a given body are 0.22 and 0.74 respectively. Then, at a given temperature, the coefficient of emission for the body is
(a) 0.4
(b) 0.04
(c) 0.96
(d) 0.22

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option b

Regards
Vikas TU
3 years ago
For a grey body
a+r+t=1 if α=0.4,r=0.6
then t=1−0.4−0.6=0
The body is opaque
////////////////////////////////////////////////////////